DIRC Operations Weekly Report

Period: Saturday 2006/08/12 - Friday 2006/08/18

Oncall: Bradley Wogsland
Fast Monitoring: Bradley Wogsland, Rolf Andreassen

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060818.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060818 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 12.07*1033 (previous week: 11.87*1033)

BaBar integrated lumi: 4.58 fb-1 (previous week: 4.56 fb-1)

Run 5 total delivery: 153.86 fb-1 Run 5 total BaBar recorded: 148.55 fb-1 including ~5 fb-1 off-peak data

PEP activities this week:

PEP activities planned next week :

Repairs in IR-4.

Foreseen work opportunities (PEP down periods):

Down until January.

BaBar activities/issues this week:

DIRC activities

The current in the bucking coil seems to really be falling according to Walt (that VSAM values rather than those from MCC are correct). Thus we have also started monitoring the B-Field inside the DIRC in case a corresponding drift appears there. None was observed before the end of the run.

Nicholas added ROM occupancy plots for each of the 6 ROMs to the fast monitoring.

See DIRC logbook for details

Backgrounds

Background summary plot

Scheduled activities

None

Pending Work

None

Manpower situation

Good.

Links